What is $1,000 invested today at 8% compounded monthly worth in 20 years?

What is $1,000.00 received every month for 5 years plus $5,000 received after all the other payments worth today if the interest rate is 8%

Please show work or how to do it on the calculator.
Answer #1
a) Future value =fv(rate,nper,pmt,pv) rate = 8%/12 = 0.00666667
= $   4,926.80 nper = 20*12 = 240
pmt = 0
pv = $ -1,000.00
b) Present value =-pv(rate,nper,pmt,fv) rate = 8%/12 = 0.00666667
= $ 52,674.49 nper = 5*12 = 60
pmt = $   1,000.00
fv = $   5,000.00